Find the LCM and HCF of the following pairs of integers and verify that LCM × HCF = product of the two numbers.
26 and 91 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Expressing 26 and 91 as product of its prime factors, we get,
26 = 2 × 13 × 1
91 = 7 × 13 × 1
Therefore, LCM (26, 91) = 2 × 7 × 13 × 1 = 182
And HCF (26, 91) = 13 Verification
Now, product of 26 and 91 = 26× 91 = 2366
And Product of LCM and HCF = 182 × 13 = 2366
Hence , LCM × HCF = product of the 26 and 91.

​Thanks